# Method-Based Reasoning for Large Language Models: Extraction, Reuse, and Continuous Improvement

### v1

@@ I Introduction
Large language models (LLMs) have achieved remarkable success in a wide range of natural language processing tasks, including text generation [1], question answering [2], and dialogue systems [3]. These models are typically built on Transformer architectures [4] [liu2025attention], which learn statistical relationships between words by training on massive corpora of text data. During inference, the model predicts the most likely next token given a sequence of known tokens, leveraging patterns learned from its training data [5].
@@ III-A1 Method Storage Tree
The tree structure is built primarily based on the problem descriptions, allowing similar problems to be grouped together. Although solutions may also be used for organization, they are discussed separately under Method Division Storage. The process begins with an initially empty tree. When a new method is introduced, the LLM is prompted to determine the appropriate location in the existing tree for insertion. Or if there are some methods, the LLM is then used to group existing methods into a tree based on problem similarity.
@@
III-A2 Method Division and Part-Wise Storage
In addition to storing complete methods, we enable fine-grained storage and manipulation of individual method components, referred to as method parts. A method typically consists of multiple procedural steps, each corresponding to a distinct subtask. By decomposing methods into parts, we support modular reuse and targeted refinement.
Formally, a method is represented as an ordered sequence of parts:
$\text{method}=\langle\text{method}_{p1},\text{method}_{p2},\dots,\text{method}_{pn}\rangle$ (5)
Each part $\text{method}_{pi}$ is associated with a specific functional role within the overall solution. The entire method is indexed by its logical characteristics, while each part can be stored, retrieved, and updated independently.
To support refinement, we define a ranking function $\text{Rank}(\cdot)$ that scores candidate method parts based on their quality or effectiveness. For a given position $i$, let $\mathcal{C}_{i}$ denote the set of candidate parts. The updated part at position $i$ is selected by:
$\text{method}_{pi}^{\text{new}}=\arg\max_{\text{candidate}\in\mathcal{C}_{i}}\text{Rank}(\text{candidate})$ (6)
The updated method sequence then becomes:
$\text{method}^{\text{new}}=\langle\text{method}_{p1},\dots,\text{method}_{pi}^{\text{new}},\dots,\text{method}_{pn}\rangle$ (7)
This part-wise update mechanism allows incremental improvements to methods by replacing only underperforming components while retaining effective ones. It also facilitates generalization, as common method parts can be reused across different tasks when subtasks overlap.
@@ VI Conclusion
Our model introduces a novel mechanism for aligning logic, not just output preferences. It integrates reinforcement-style feedback from users to rank and refine the extracted methods over time. By associating each method with an abstract representation of its corresponding problem and solution, the system supports generalization to new tasks that share similar logical structures, even if their surface forms differ significantly.
@@
The proposed system bridges the gap between symbolic planning and neural generation. It enables LLMs to go beyond token-level prediction by equipping them with reusable and interpretable procedural knowledge. This allows for increased consistency, explainability, and adaptability in dynamic or complex interaction settings.

### v2

@@ I Introduction
Large language models (LLMs) have achieved remarkable success in a wide range of natural language processing tasks, including text generation [1], question answering [2], and dialogue systems [3]. These models are typically built on Transformer architectures [4] [5], which learn statistical relationships between words by training on massive corpora of text data. During inference, the model predicts the most likely next token given a sequence of known tokens, leveraging patterns learned from its training data [6].
@@ III-A1 Method Storage Tree
The tree structure is built primarily based on the problem descriptions, allowing similar problems to be grouped together. Although solutions may also be used for organization, they are discussed separately under section of III-A2 and III-A3. The process begins with an initially empty tree. When a new method is introduced, the LLM is prompted to determine the appropriate location in the existing tree for insertion. Or if there are some methods, the LLM is then used to group existing methods into a tree based on problem similarity.
@@
III-A2 Methods for Improving Other Methods
In our model, a method (or solution) is typically associated with solving a specific user problem. However, there are scenarios where a method itself may become the subject of another method—especially when the original method fails, produces suboptimal results, or requires refinement.
For instance, if a method produces an incorrect or incomplete solution, another method can be invoked to validate, revise, or enhance it. One common example is the use of Chain-of-Thought (CoT) prompting to transform a shallow method into a more structured, step-by-step reasoning process. This type of meta-method can be seen as a higher-level strategy that operates on other methods rather than on raw input problems.
We formalize this idea as follows. Let $\mathcal{M}$ denote the set of all methods, and let $\text{Apply}(m,q)$ denote the application of method $m$ to a query $q$ or a method $m$ to produce an output. Then, for a method $m_{1}\in\mathcal{M}$ and an input $q$, we define its output as:
$o_{1}=\text{Apply}(m_{1},q)$ (5)
If the output $o_{1}$ is deemed unreliable, a secondary method $m_{2}$ can be applied to the original method $m_{1}$ (with knowledge of $q$) to improve its behavior:
$o_{2}=\text{Apply}(m_{2},m_{1},q)$ (6)
Here, $m_{2}$ is acting on $m_{1}$ as its input context—serving as a higher-level reasoning strategy. It may, for instance, prompt the system to re-express $m_{1}$ using CoT or add a validation step before accepting $m_{1}$’s result.
Thus, a hierarchical structure emerges, where:
- Leaf nodes represent base-level methods that address user-level queries.
- Internal nodes represent higher-level methods that operate on other methods, refining or validating their behavior.
We define a meta-method relationship as follows:
$m_{2}\Rightarrow m_{1}\quad\text{(i.e., $m_{2}$ refines or validates $m_{1}$)}$ (7)
Thus, a hierarchical structure enables the system to evolve complex strategies by composing and improving methods over time. The architecture supports recursive reasoning, layered validation, and the ability to adapt existing procedures for improved performance and reliability.
III-A3 Solution Division and Part-Wise Storage
In addition to storing complete solutions, we enable fine-grained storage and manipulation of individual solution components, referred to as solution parts. A solution typically consists of multiple procedural steps, each corresponding to a distinct subtask. By decomposing solutions into parts, we support modular reuse and targeted refinement.
Formally, a solution is represented as an ordered sequence of parts:
$\text{solution}=\langle\text{solution}_{p1},\text{solution}_{p2},\dots,\text{solution}_{pn}\rangle$ (8)
Each part $\text{solution}_{pi}$ is associated with a specific functional role within the overall solution. The entire solution is indexed by its logical characteristics, while each part can be stored, retrieved, and updated independently.
To support refinement, we define a ranking function $\text{Rank}(\cdot)$ that scores candidate solution parts based on their quality or effectiveness. For a given position $i$, let $\mathcal{C}_{i}$ denote the set of candidate parts. The updated part at position $i$ is selected by:
$\text{solution}_{pi}^{\text{new}}=\arg\max_{\text{candidate}\in\mathcal{C}_{i}}\text{Rank}(\text{candidate})$ (9)
The updated solution sequence then becomes:
$\text{solution}^{\text{new}}=\langle\text{solution}_{p1},\dots,\text{solution}_{pi}^{\text{new}},\dots,\text{solution}_{pn}\rangle$ (10)
This part-wise update mechanism allows incremental improvements to solutions by replacing only underperforming components while retaining effective ones. It also facilitates generalization, as common solution parts can be reused across different tasks when subtasks overlap.
@@ V Conclusion
Our model introduces a mechanism for aligning logic, not just output preferences. It integrates reinforcement-style feedback from users to rank and refine the extracted methods over time. By associating each method with an abstract representation of its corresponding problem and solution, the system supports generalization to new tasks that share similar logical structures, even if their surface forms differ significantly.